# Comprehensive Review of Deep Unfolding Techniques for Next-Generation Wireless Communication Systems

### v2

@@ XII-B1 Coexistence of Deep Unfolding Models, Conventional Solvers, and Black-Box Models
XII-B2 Deep Unfolding for Multi-Objective Learning
Most existing deep unfolding models are developed for single-objective designs, such as maximizing throughput, SINR, or minimizing BER. However, in many scenarios, system designs involve multiple objectives. Traditional approaches often formulate these as constrained problems, such as maximizing energy efficiency under a QoS constraint (e.g., throughput or SINR). In ISAC designs, weighted sums of communication and sensing performance metrics are used when the goal is to balance both functions without prioritizing one over the other. The complexity grows in multi-static ISAC systems, where distributed sensing receivers have different performance metrics depending on their functions, such as SCNR, mutual information, Cramér-Rao lower bound, or probability of detection.
@@ XII-B2 Deep Unfolding for Multi-Objective Learning
XII-B3 Distributed Deep Unfolding Architectures
While deep unfolding models provide advantages in terms of complexity due to faster convergence compared to conventional iterative algorithms, they can still face high complexity issues due to challenge C2. This limits their applicability on platforms with constrained computational resources, such as edge devices or distributed access points (APs). Future wireless networks will advance beyond simple communications, evolving into large-scale, distributed systems that integrate advanced sensing, communications, and computing capabilities. However, most existing AI models, both data-driven and model-driven, were designed for centralized scenarios, where they are deployed at a central node with substantial computing power. These models are not well-suited for distributed scenarios, highlighting the need for the development of appropriate distributed unfolding architectures.
@@ XII-B3 Distributed Deep Unfolding Architectures
XII-B4 Deep Unfolding to Enhance Black-box Solution
@@
As its name suggests, a black-box ML model operates with a black-box nature, making it unexplainable and resistant to re-structuring beyond tuning parameters like the number of layers or nodes. The performance and efficiency of such models are heavily dependent on the quality of input data. Deep unfolding can contribute by performing pre-processing steps to improve the input data quality. One example is in multi-modal sensing-assisted communication tasks, where data from distributed cameras, LiDAR, and radar sensors are used to enhance communication performance, such as in beam tracking and blockage prediction. However, significant challenges arise in collecting, transferring, and processing large sensory data. Specifically, high-resolution camera images generate considerable traffic overhead and latency, impacting real-time communication processing.
Deep unfolding models can be particularly useful in pre-processing such images. In image processing, deep unfolding has been successfully applied to tasks such as denoising, super-resolution, and image reconstruction, allowing for high-quality outcomes with fewer iterations than conventional algorithms. For example, unfolded algorithms based on the ISTA for sparse coding have been shown to significantly accelerate image reconstruction while maintaining high-quality results [293, 294, 295]. This makes deep unfolding a valuable tool for handling large-scale sensory data in real-time communication systems.
XIII Conclusions
@@
The integration of traditional algorithmic principles with DL techniques has led to the development of deep unfolding, a method that overcomes the limitations associated with both conventional algorithms and standalone DL models in wireless communication. This article offers a comprehensive survey of deep unfolding techniques, examining their application across various facets of the physical layer in wireless communication systems. Key applications, such as signal detection, channel estimation, precoder design, sensing, communication decoding for error-correcting codes, power allocation, and security, are explored, with an emphasis on technologies anticipated for beyond-5G networks. The survey highlights the strengths of deep unfolding in addressing complex challenges, particularly in achieving low-latency, high-accuracy performance in dynamic environments. We conclude by discussing the key challenges in developing deep unfolding models, including issues related to scalability, computational efficiency, and robustness across diverse scenarios. Finally, we propose potential advancements to enhance the adaptability and resilience of deep unfolding, with the goal of solidifying its role as a robust and versatile technique for next-generation wireless communications.

### v3

@@ X-B1 Coexistence of Deep Unfolding Models, Conventional Solvers, and Black-Box Models
X-B2 Deep Unfolding for Multi-Objective Learning
Most existing deep unfolding models are developed for single-objective designs, such as maximizing throughput, SINR, or minimizing BER. However, in many scenarios, system designs involve multiple objectives. Traditional approaches often formulate these as constrained problems, such as maximizing energy efficiency under a QoS constraint (e.g., throughput or SINR). In ISAC designs, weighted sums of communication and sensing performance metrics are used when the goal is to balance both functions without prioritizing one over the other. The complexity grows in multi-static ISAC systems, where distributed sensing receivers have different performance metrics depending on their functions, such as signal clutter noise ratio (SCNR), mutual information, CRB, or probability of detection.
@@ X-B2 Deep Unfolding for Multi-Objective Learning
X-B3 Distributed Deep Unfolding Architectures
While deep unfolding models provide advantages in terms of complexity due to faster convergence compared to conventional iterative algorithms, they can still face high complexity issues due to challenge C2. This limits their applicability on platforms with constrained computational resources, such as edge devices or distributed access points (APs). Future wireless networks will advance beyond simple communications, evolving into large-scale, distributed systems such as cell-free massive MIMO that integrate advanced sensing, communications, and computing capabilities. However, most existing AI models, both data-driven and model-driven, were designed for centralized scenarios, where they are deployed at a central node with substantial computing power. These models are not well-suited for distributed scenarios, highlighting the need for appropriate distributed unfolding architectures.
@@ X-B3 Distributed Deep Unfolding Architectures
X-B4 Deep Unfolding to Enhance Black-box Solution
@@
As its name suggests, a black-box ML model operates with a black-box nature, making it unexplainable and resistant to re-structuring beyond tuning parameters like the number of layers or nodes. The performance and efficiency of such models are heavily dependent on the quality of input data. Deep unfolding can contribute by performing pre-processing steps to improve the input data quality. One example is in multi-modal sensing-assisted communication tasks, where data from distributed cameras, LiDAR, and radar sensors are used to enhance communication performance, such as in beam tracking and blockage prediction. However, significant challenges arise in collecting, transferring, and processing large sensory data. Specifically, high-resolution camera images generate considerable traffic overhead and latency, impacting real-time communication processing. Deep unfolding models can be particularly useful in pre-processing such high resolution images. In image processing, deep unfolding has been successfully applied to tasks such as denoising, super-resolution, and image reconstruction, allowing for high-quality outcomes with fewer iterations than conventional algorithms. For example, unfolded algorithms based on ISTA for sparse coding have been shown to significantly accelerate image reconstruction while maintaining high-quality results [235, 236, 237]. This makes deep unfolding a valuable tool for handling large-scale sensory data in real-time communication systems.
XI Conclusions
@@
The integration of classical algorithmic principles with DL techniques has led to the development of deep unfolding, a method that overcomes the limitations associated with both classical algorithms and standalone DL models in wireless communication. This article offers a comprehensive survey of deep unfolding techniques, examining their application across various facets of the physical layer in wireless communication systems. Key applications, such as signal detection, channel estimation, precoder design, ISAC, decoding for error-correcting codes, power allocation, and physical layer security, are explored, with an emphasis on technologies anticipated for beyond-5G networks. The comparisons of the unfolded networks have also been given, which reflect the performance gains over the benchmark iterative algorithms. Also, we have analyzed the advantages and limitations of the design of various unfolding models. Moreover, we have explored the practical feasibility of deep unfolding networks across key application areas of wireless communications. The survey highlights the strengths of deep unfolding in addressing complex challenges, particularly in achieving low-latency, high-accuracy performance in dynamic environments. We conclude by discussing the key challenges in developing deep unfolding models, including issues related to scalability, computational efficiency, and robustness across diverse scenarios. Finally, this paper presents potential advancements to enhance the adaptability and resilience of deep unfolding, with the goal of solidifying its role as a robust and versatile technique for next-generation wireless communications.